A motorcycle daredevil is planning a stunt to perform at a county fair. A ramp must be built to give him a 20% grade, or slope. If the vertical height at the end of the ramp must be 11 ft to assure that the stunt is a success, what must be the length of the horizontal run?

Answer by Nate(3500)

Slope: 20% of 90 = (2/10)90 = 18
Sin(18 degrees) = 11/x
x = 11/Sin(18pi/180) = 36 ft